# Searching for Particle Dark Matter with eROSITA Early Data

### v2

@@ III.5 Sterile Neutrino Dark Matter Constraints
In figure 6, we show sterile neutrino DM constraint derived from eFEDS, projected future eROSITA sensitivities, and all the existing constraints. In eROSITA energy space, relevant previous X-ray surveys include: Chandra Horiuchi et al. (2014); Sicilian et al. (2020), XMM-Newton Watson et al. (2006); Dessert et al. (2020); Foster et al. (2021), NuSTAR Neronov et al. (2016); Perez et al. (2017); Ng et al. (2019); Roach et al. (2020, 2023), and Suzaku Loewenstein et al. (2009). The limit obtained by eFEDS in this work is the strongest X-ray limit in - (see also figure 7 in section III.6). This is due to the long exposure time of eFEDS, which, after co-adding up all 7 TMs, gives of exposure time. Our limits are roughly consistent with previously eROSTIA projections Dekker et al. (2021), and SRG sterile neutrino search with correlation function Barinov (2023)
Sterile neutrino DM can be produced through mixing with active neutrinos in the early universe with Shi-Fuller mechanism Shi and Fuller (1999). This mechanism proposed that in the presence of large lepton asymmetry, even a small mixing angle could be enhanced by extra matter potential, therefore producing the amount of DM needed. In the special case of zero lepton asymmetry, and thus the production is non-resonant, it is known as Dodelson-Widrow mechanism Dodelson and Widrow (1994). Assuming sterile neutrino DM is produced by Shi-Fuller mechanism, one also needs to consider constraint from big bang nuclear synthesis (BBN), on the amount of lepton asymmetry in the universe Canetti et al. (2013a). We show the lower bound on Boyarsky et al. (2009); Laine and Shaposhnikov (2008); Dolgov and Hansen (2002), which is calculated from sterile-dm Venumadhav et al. (2016), given the lepton asymmetry consideration from BBN, .
@@ Acknowledgments
The works of CF, KCYN and QSL are supported by Croucher foundation, RGC grants (24302721, 14305822, 14308023), and NSFC/GRC grant (N_CUHK456/22).
@@ Appendix A Observed Data and Fitted Model of Each Individual Telescope Module
In the main text, we report the fitting procedure for our blank sky and DM model. We only show the best fit result for TM1 as an example in figure 2. Here, we report the best fit results for all 7 TMs in figure A.1 and all frozen parameter values in table 4. There are 171 degrees of freedom for TMs not affected by light leak, and 170 d.o.f. for TM5 and TM7.
@@
Origin | Model name | Model Param. | Telescope Module | |||||||
1 | 2 | 3 | 4 | 5 | 6 | 7 | ||||
Astrophysical | apec | abund | 1 | |||||||
0 | |||||||||
plaw |
1.46 | ||||||||
tbabs | ||||||||||
Instrument | gaussian | line energy | 9.57 | 9.63 | 9.57 | 9.68 | 9.66 | 9.50 | 9.63 | |
8.85 | 8.85 | 8.90 | 8.85 | 8.90 | 8.85 | 8.90 | ||||
8.20 | 8.20 | 8.23 | 8.20 | 8.30 | 8.20 | 8.20 | ||||
8.59 | 8.62 | 8.61 | 8.62 | 8.64 | 8.59 | 8.62 | ||||
8.04 | 8.10 | 8.03 | 8.06 | 8.10 | 8.05 | 8.05 | ||||
7.44 | 7.48 | 7.46 | 7.47 | 7.50 | 7.46 | 7.48 | ||||
7.03 | 7.03 | 7.04 | 7.06 | 7.04 | 7.00 | 7.07 | ||||
6.91 | 6.85 | 6.91 | 6.89 | 7.01 | 6.93 | 6.95 | ||||
6.38 | 6.40 | 6.40 | 6.40 | 6.42 | 6.39 | 6.41 | ||||
5.88 | 5.87 | 5.86 | 5.95 | 5.87 | 5.88 | 5.91 | ||||
5.38 | 5.35 | 5.39 | 5.42 | 5.49 | 5.55 | 5.39 | ||||
4.52 | 4.52 | 4.50 | 4.50 | 4.52 | 4.52 | 4.50 | ||||
3.65 | 3.71 | 3.68 | 3.73 | 3.65 | 3.65 | 3.73 | ||||
1.48 | 1.49 | 1.49 | 1.49 | 1.49 | 1.48 | 1.49 | ||||
0.0001 | |||||||||
exponential modification | -2.12 | -2.13 | -1.66 | -1.48 | -2.28 | -2.06 | -1.75 | |||
0.05 | |||||||||
broken power law, 2 break energies | 0.06 | 0.07 | 0.08 | 0.11 | 0.14 | 0.07 | 0.13 | |||
3.62 | 6.56 | 4.51 | 8.99 | 9.24 | 6.26 | 96.9 | |||
0.13 | 0.28 | 0.19 | 3.3 | -2.5 | 0.29 | 0.03 | ||||
9.13 | 9.16 | 9.64 | 9.68 | 9.56 | 9.31 | 9.5 | |||
6.5 | -2.5 | 4.59 | 9.5 | 0.41 | 9.5 | 0.23 | ||||
power law | 2.18 | 2.22 | 2.12 | 2.84 | 5.26 | 1.95 | 2.74 | |||
power law | 6.0 | 6.0 | 6.0 | 6.0 | 6.0 | 9.5 | 6.0 | |||
Free | Free | Free | Free | 0.0 | Free | 0.0 |

### v3

@@ III.5 Sterile Neutrino Dark Matter Constraints
In figure 6, we show sterile neutrino DM constraint derived from eFEDS, projected future eROSITA sensitivities, and all the existing constraints. In eROSITA energy space, relevant previous X-ray surveys include: Chandra Horiuchi et al. (2014); Sicilian et al. (2020), XMM-Newton Watson et al. (2006); Dessert et al. (2020); Foster et al. (2021), NuSTAR Neronov et al. (2016); Perez et al. (2017); Ng et al. (2019); Roach et al. (2020, 2023), and Suzaku Loewenstein et al. (2009). The limit obtained by eFEDS in this work is the strongest X-ray limit in - (see also figure 7 in section III.6). This is due to the long exposure time of eFEDS, which, after co-adding up all 7 TMs, gives of exposure time. Our limits are roughly consistent with previously eROSITA projections Dekker et al. (2021), and SRG sterile neutrino search with correlation function Barinov (2023)
Sterile neutrino DM can be produced through mixing with active neutrinos in the early universe with Shi-Fuller mechanism Shi and Fuller (1999). This mechanism proposed that in the presence of large lepton asymmetry, even a small mixing angle could be enhanced by extra matter potential, therefore producing the amount of DM needed. In the special case of zero lepton asymmetry, and thus the production is non-resonant, it is known as Dodelson-Widrow mechanism Dodelson and Widrow (1994). Assuming sterile neutrino DM is produced by Shi-Fuller mechanism, one also needs to consider constraint from big bang nucleosynthesis (BBN), on the amount of lepton asymmetry in the universe Canetti et al. (2013a). We show the lower bound on Boyarsky et al. (2009); Laine and Shaposhnikov (2008); Dolgov and Hansen (2002), which is calculated from sterile-dm Venumadhav et al. (2016), given the lepton asymmetry consideration from BBN, .
@@ Acknowledgments
The works of CF, KCYN and QSL are supported by KCYN is supported by Croucher foundation, RGC grants (24302721, 14305822, 14308023), NSFC/RGC joint research scheme (N CUHK456/22), and NSFC grant 12322517. This work is supported in part by the National Natural Science Foundation of China under grant No. 12342502.
@@ Appendix A Observed Data and Fitted Model of Each Individual Telescope Module
In the main text, we report the fitting procedure for our blank sky and DM model. We only show the best fit result for TM1 as an example in figure 2. Here, we report the best fit results for all 7 TMs in figure A.1 and all best fit parameter values in table 3. There are 145 degrees of freedom for TMs not affected by light leak, and 147 d.o.f. for TM5 and TM7.
@@
| Origin | Model name | Model Param. (unit) | Telescope Module | Freeze Status (Range) | |||||||||
| 1 | 2 | 3 | 4 | 5 | 6 | 7 | |||||||
| Astrophysical | apec | Abund (Solar) | 1 | 1 | 0.99 | 1.00 | 1.00 | 0.99 | 1.00 | Free | |||
0 | Frozen | |||||||||||
0.67 | 0.76 | 0.70 | 0.76 | 0.71 | 0.66 | 0.86 | Free (0.5, 1) | |||||
| Normalization | 0.092 | 0.065 | 0.075 | 0.064 | 0.047 | 0.061 | 0.041 | Free | |||||
| power law |
1.51 | 1.44 | 1.50 | 1.46 | 1.58 | 1.47 | 1.50 | Free (-3, 10) | ||||
| Normalization | 0.37 | 0.34 | 0.37 | 0.35 | 0.34 | 0.29 | 0.36 | Free | |||||
| tbabs |
4.4 | 4.4 | 4.4 | 4.4 | 4.4 | 4.4 | 4.4 | Free | ||||
| Instrument | gaussian |
9.80 | 9.80 | 9.57 | 9.80 | 9.80 | 9.80 | 9.80 | Free (9.45, 9.8) | |||
| Normalization | 1.78e-18 | 1.78e-18 | 0.00 | 1.59e-18 | 5.05e-18 | 7.89e-19 | 2.34e-18 | Free | |||||
8.85 | 8.85 | 8.85 | 8.85 | 9.10 | 8.85 | 8.85 | Free (8.8, 9.15) | |||||
| Normalization | 4.19e-19 | 4.34e-07 | 7.20e-19 | 5.17e-04 | 7.89e-18 | 8.63e-19 | 9.08e-04 | Free | |||||
8.30 | 8.30 | 8.23 | 8.30 | 8.30 | 8.20 | 8.23 | Free (8.15, 8.35) | |||||
| Normalization | 1.90e-05 | 2.99e-18 | 8.27e-04 | 4.96e-19 | 1.75e-04 | 1.55e-18 | 1.39e-03 | Free | |||||
8.60 | 8.63 | 8.62 | 8.62 | 8.66 | 8.59 | 8.63 | Free (7.45, 8.75) | |||||
| Normalization | 5.26e-03 | 5.46e-03 | 4.22e-03 | 4.35e-03 | 4.57e-03 | 5.48e-03 | 6.99e-03 | Free | |||||
8.04 | 8.09 | 8.07 | 8.10 | 8.10 | 8.07 | 8.08 | Free (7.8, 8.15) | |||||
| Normalization | 2.40e-03 | 1.61e-03 | 2.51e-03 | 2.76e-03 | 1.95e-03 | 2.76e-03 | 1.57e-03 | Free | |||||
7.47 | 7.49 | 7.49 | 7.47 | 7.50 | 7.47 | 7.48 | Free (7.2, 7.6) | |||||
| Normalization | 3.45e-03 | 3.51e-03 | 4.91e-03 | 4.33e-03 | 4.48e-03 | 4.54e-03 | 4.92e-03 | Free | |||||
7.01 | 6.97 | 7.05 | 7.02 | 7.05 | 7.04 | 6.95 | Free (6.9, 7.2) | |||||
| Normalization | 2.75e-03 | 2.69e-03 | 2.11e-03 | 1.26e-03 | 1.26e-03 | 1.70e-03 | 8.12e-04 | Free | |||||
6.86 | 7.05 | 6.91 | 6.94 | 6.97 | 6.94 | 7.01 | Free (6.8, 7.1) | |||||
| Normalization | 9.23e-04 | 1.37e-18 | 9.54e-04 | 6.49e-04 | 1.20e-03 | 1.12e-03 | 1.36e-03 | Free | |||||
6.39 | 6.41 | 6.41 | 6.41 | 6.42 | 6.40 | 6.41 | Free (6.2, 6.6) | |||||
| Normalization | 0.013 | 0.011 | 0.012 | 0.012 | 0.012 | 0.011 | 0.012 | Free | |||||
5.90 | 5.94 | 5.93 | 5.92 | 5.95 | 5.90 | 5.92 | Free (5.7, 6) | |||||
| Normalization | 1.76e-03 | 1.76e-03 | 1.46e-03 | 7.50e-04 | 9.91e-04 | 3.12e-04 | 1.64e-03 | Free | |||||
5.44 | 5.55 | 5.40 | 5.42 | 5.55 | 5.55 | 5.47 | Free (5.3, 5.6) | |||||
| Normalization | 8.67e-04 | 5.10e-19 | 2.67e-04 | 6.34e-04 | 1.25e-17 | 2.70e-19 | 8.99e-04 | Free | |||||
4.50 | 4.50 | 4.52 | 4.52 | 4.52 | 4.52 | 4.52 | Free (4.49, 4.53) | |||||
| Normalization | 3.26e-04 | 8.63e-04 | 1.61e-18 | 1.13e-04 | 6.49e-04 | 2.29e-18 | 8.60e-04 | Free | |||||
3.67 | 3.68 | 3.68 | 3.66 | 3.68 | 3.67 | 3.72 | Free (3.6, 3.8) | |||||
| Normalization | 1.08e-03 | 1.63e-03 | 1.76e-03 | 1.10e-03 | 1.20e-03 | 1.08e-03 | 1.07e-03 | Free | |||||
1.48 | 1.49 | 1.49 | 1.49 | 1.48 | 1.48 | 1.49 | Free (1.47, 1.5) | |||||
| Normalization | 4.94e-03 | 5.23e-03 | 4.66e-03 | 5.48e-03 | 3.19e-03 | 6.22e-03 | 2.99e-03 | Free | |||||
0.0001 | Frozen | |||||||||||
| exponential modification | Amplitude | -1.67e-10 | -1.40e-10 | -9.69e-09 | -1.02e-07 | -7.45e-11 | 2.46e-10 | -1.27e-08 | Free (-2, 2) | ||||
-2.10 | -2.12 | -1.68 | -1.49 | -2.25 | -2.03 | -1.75 | Free (-4, 4) | |||||
0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | 0.05 | Free | |||||
| broken power law, 2 break energies | 0.06 | 0.07 | 0.09 | 0.11 | 0.18 | 0.07 | 0.14 | Free (-3, 10) | |||||
3.96 | 5.26 | 4.71 | 8.89 | 8.92 | 6.23 | 96.90 | Free | |||||
| 0.14 | 0.19 | 0.17 | -2.50 | -2.50 | 0.28 | 0.03 | Free (-3, 10) | ||||||
8.89 | 9.35 | 18.08 | 12.79 | 10.84 | 8.87 | 9.50 | Free | |||||
| 7.71 | -2.31 | 9.15 | 2.00 | -1.54 | 9.50 | 0.23 | Free (-3, 10) | ||||||
| Normalization | 0.20 | 0.19 | 0.20 | 0.21 | 0.27 | 0.20 | 0.24 | Free | |||||
| power law |
1.24 | 2.05 | 0.80 | 2.34 | 5.85 | 1.86 | 1.73 | Free (-3, 10) | ||||
| Normalization | 3.77e-03 | 0.05 | 2.65e-03 | 0.01 | 0.02 | 0.05 | 0.06 | Free | |||||
| power law |
2.78 | 6.00 | 2.35 | 8.83 | 6.00 | 1.21 | 6.00 |
0.03 | 0.00 | 0.02 | 9.68e-04 | 0.00 | 3.22e-05 | 0.00 |
Appendix B The Impact of Using Different CXB Power Law Indices on DM Constraint
We investigate how different models of CXB could impact our DM constraint result. Since CXB is modeled by a power law and we allow all normalization to float freely in our DM search procedure, the only choice in CXB parameter is the index of power law. In figure B.1 we show 4 constraint results with different choices of power law indices: (1) our fiducial result, where power law index is unfrozen (free) in initial spectral fitting and DM search, with the best values of the 7 TMs falling between 1.46-1.58; (2) power law index frozen at 1.42, which is reported in XMM-Newton papers and suitable for hard X-ray sky Bulbul et al. (2012); (3) frozen at 1.5 which is close to the average of our best fit null models for 7 individual TMs; (4) frozen at 1.7 which is reported in eROSITA diffuse sky study and suitable for soft X-ray sky Ponti et al. (2023). As shown in the bottom panel of figure B.1, we find that the difference choices of CXB power law model indices have minor impact on our DM constraint result. The ratios of DM constraints from different CXB models to our fiducial result are within factor of 3, except in narrow ranges of energy where the fit is problematic for all models.